Outline the steps for preparing 0.05M sodium hydroxide solution in a 500ml volumetric flask.

You are provided the following:
-a weighing balance
-Sodium hydroxide pellets
-distilled water
-a weighing bottle
-100ml beaker
-500ml volumetric flask
-stirring rod
1
Expert's answer
2018-05-29T06:34:18-0400
С(NaOH)=n(NaOH)/V(NaOH)
n(NaOH)=С(NaOH)*V(NaOH)= 0.05M*0.5L= 0.025mol
m(NaOH)=n(NaOH)*M(NaOH)=0.025 mol*40g/mol = 1g
Starting with anhydrous, solid NaOH weigh out 1.00 g and carefully add it to an empty 500.00 mL volumetric flask. Add about 400 mL of distilled H2O, stopper, shake to dissolve. The bottle will become warm to hot to touch as NaOH has an exothermic heat of solution. Let it cool to RT and add distilled water carefully until the bottom of the meniscus touches the 500.00 mL line. For greatest accuracy you actually need a temperature controlled room set to 20.0 degrees because all volumetric glassware is only accurate to its stated error at that temp.